# Multi-Token Joint Speculative Decoding for Accelerating Large Language Model Inference

### v1

@@ Abstract
Transformer-based Large language models (LLMs) have demonstrated their power in various tasks, but their inference incurs significant time and energy costs. To accelerate LLM inference, speculative decoding uses a smaller model to propose one sequence of tokens, which are subsequently validated in batch by the target large model. Compared with autoregressive decoding, speculative decoding generates the same number of tokens with fewer runs of the large model, hence accelerating the overall inference by $1$-2×2$\times$2 ×. However, greedy decoding is not the optimal decoding algorithm in terms of output perplexity, which is a direct measurement of the effectiveness of a decoding algorithm. An algorithm that has better output perplexity and even better efficiency than speculative decoding can be more useful in practice. To achieve this seemingly contradictory goal, we first introduce multi-token joint greedy decoding (MJGD), which greedily generates multiple tokens at each step based on their joint perplexity. We show that it leads to better perplexity for the whole output. But the computation cost of MJGD is infeasible in practice. So we further propose multi-token joint speculative decoding (MJSD), which approximates and accelerates the MJGD from two aspects: it approximates the joint distribution of the large model with that of a small model, and uses a verification step to guarantee the accuracy of approximation; then it uses beam decoding to accelerate the sequence generation from the joint distribution. Compared with vanilla speculative decoding, MJSD has two advantages: (1) it is an approximation of MJGD, thus achieving better output perplexity; (2) verification with joint likelihood allows it to accept the longest prefix sub-sequence of the draft tokens with valid perplexity, leading to better efficiency. In addition, we analyze energy consumption during LLM inference and provide theoretical and empirical evidence that, surprisingly, MJSD reduces energy consumption even when the number of operations increases. Experiment results show that our approach enhances perplexity by $21.2\%$ than greedy decoding. Moreover, MJSD achieves 2.21×2.21$\times$2.21 × speed-up and 2.84×2.84$\times$2.84 × less energy consumption than greedy decoding, and 1.49×1.49$\times$1.49 × speed-up and 1.62×1.62$\times$1.62 × less energy consumption than vanilla speculative decoding.
@@ 1 Introduction
Large Language Models (LLMs) have exhibited remarkable performance across real-world tasks spanning text and image domains [4, 7, 33, 34]. However, the substantial parameter size of these models leads to high computational cost during inference [27]. For example, ChatGPT is estimated to have an energy demand of 564 MWh per day [8]. Compounding this issue, the autoregressive generation of $K$ tokens necessitates calling the model $K$ times, involving the repeated loading of weight matrices and intermediate results from GPU global memory to computing units. As a result, LLM inference is often hampered by constraints related to memory bandwidth and communication [20], leading to diminished hardware utilization, heightened inference latency, and elevated energy cost.
@@
To tackle this challenge, researchers have delved into non-autoregressive decoding approaches. Earlier methods [11, 12, 13] aimed at reducing inference latency by concurrently generating multiple tokens. But these methods usually require task-dependent techniques and information to achieve comparable performance to that of autoregressive decoding [17, 36]. More recently, speculative decoding has emerged [20, 5, 17, 32], exploiting the observation that, even when a small model is ten times smaller than a large model, most of the small model’s prediction aligns well with that of a large model. So it leverages a smaller auxiliary model to draft a few future tokens autoregressively, which are subsequently validated in parallel by the larger model. As the smaller model operates significantly faster and parallel token verification incurs a similar time cost as generating a single token, speculative decoding achieves an overall speed-up of $1$-2×2$\times$2 ×. Meanwhile, it always has an identical sampling distribution as greedy decoding.
However, there are some limitations of the existing studies on speculative decoding. First, they aim to have an identical sampling distribution as greedy sampling. But the goal of decoding algorithms is to generate sequences with optimal perplexity, which is defined as the exponentiated average negative log-likelihood of all tokens. An algorithm that does not follow the same sampling distribution as greedy decoding, but yields output with better perplexity could be more useful in the real world. Second, in these algorithms, if a draft token is rejected, all the draft tokens after it are discarded. But the overall draft sequences may have good quality despite there being a few low-quality tokens. Rejecting these sequences will shorten the average acceptance length for each iteration, lowering the decoding efficiency. Third, existing research does not investigate the impact of speculative decoding on inference energy consumption. While speculative decoding increases the number of FLOPs during inference, it simultaneously reduces the overall inference time, communication, and memory operations, which are crucial factors in determining energy consumption [1, 6]. Thus, it remains unclear if speculative decoding increases or decreases the energy cost of inference.
So we aim to design a new algorithm that is both more efficient and more effective than greedy decoding and vanilla speculative decoding. We first introduce multi-token joint greedy decoding (MJGD), which generates multiple tokens at each step greedily based on their joint likelihood. We empirically show that it has better overall perplexity than greedy decoding does. However, the computational cost of MJGD is infeasible in practice. So we further propose multi-token joint speculative decoding (MJSD), which approximates and accelerates MJGD from two aspects: (1) it approximates the joint distribution of the large model with that of a small model by generating draft tokens with the small model and validating them with the large model; (2) it uses beam decoding to accelerate the sequence generation from the joint distribution of the small model. Compared with vanilla speculative decoding, MJSD has two advantages: (1) it is more effective, as it approximates MJGD, which has better output perplexity than greedy decoding does; (2) it is more efficient because it uses joint likelihood as the verification criterion, which allows it to accept the longest prefix sub-sequence of the draft tokens with valid perplexity. In addition, we analyze the energy consumption of LLM inference. We give theoretical and empirical evidence that, despite that MJSD and other speculative decoding algorithms increase the number of FLOPs needed during LLM inference, they reduce the overall energy consumption by reducing the overhead induced by accessing GPU global memory.
We evaluate our method on text generation tasks with Llama-2 [34] and OPT models [38]. Experiment results show that our approach enhances perplexity by $21.2\%$ than the baselines with identical sampling distributions as greedy decoding. In addition, MJSD achieves 2.21×2.21$\times$2.21 × speed-up and 2.84×2.84$\times$2.84 × smaller energy consumption than greedy decoding, and 1.49×1.49$\times$1.49 × speed-up and 1.62×1.62$\times$1.62 × smaller energy consumption than vanilla speculative decoding. Our code is open-sourced [^1] https://anonymous.4open.science/r/LLMSpeculativeSampling-EE52.
@@ 2.1 Decodings of LLMs
Let $p$ denote the distribution defined by LLM model $M_{p}$. Given an input prefix $prefix$, the optimal decoding algorithm is to generate a sequence of $N$ tokens with maximum likelihood $p(x_{1:N}|prefix)$. The likelihood of the sequence is directly linked to perplexity of the sequence, which is the exponentiated average negative log-likelihood of all tokens. Based on autoregressive decomposition $p(x_{1:N}|prefix)=\prod_{t=1}^{N}p(x_{t}|x_{1:t-1},prefix)$, the perplexity is defined as:
@@
Greedy Decoding.
Greedy decoding samples the next single token $x_{t}$ based on $p(\cdot|x_{1:t-1},prefix)$. According to the autoregressive decomposition, greedy sampling leads to the desired joint distribution for the sequence. In practice, however, the objective of decoding algorithms is to generate sequences with high likelihood. Instead of sampling from the joint distribution, a warping operation $\mathcal{T}$ usually is applied to boost the high probability region, which transforms the original distribution $p$ to $\tilde{p}=\mathcal{T}\circ p$. A very popular warping operation is $argmax$, which only selects the sample with the highest probability. Other possible warping operations include top-k. For most of the warping operations, autoregressive decomposition no longer holds after the warping. For example, $\arg\max_{x_{1:N}}{p}(x_{1:N}|prefix)\neq\prod_{t=1}^{N}\arg\max_{x_{t}}{p}(x_{t}|x_{1:t-1},prefix)$.
Beam Decoding.
Beam decoding aims to do a better job than greedy decoding. For each position $t$ ($1\leq t\leq N$), it maintains $M>1$ candidate sequences, which are also called beams. Still take $argmax$ warping operator as an example and assume we have already kept the $M$ sequences $\mathcal{I}_{t-1}=\{x_{1:t-1}^{(1)},\ldots,x_{1:t-1}^{(M)}\}$ at position $t-1$ with highest likelihood, $M$ tokens are then sampled from $p(\cdot|x_{1:t-1}^{(m)},prefix)$ for each sequence $m$ and the top $M$ sequences with the highest likelihood $p(x_{1:t}|prefix)$ will be kept. It is widely recognized that beam decoding is a closer approximation to $\tilde{p}(x_{1:N}|prefix)$ than greedy decoding [18], with few exceptions under special construction. The computation for each beam can be parallelized, thus its run time is about the same as greedy decoding. The fact that beam decoding is more effective than greedy decoding also suggests that the output perplexity of greedy decoding still has room for improvement.
@@ 2.2 Vanilla Speculative Decoding
Besides effectiveness, speculative decoding is proposed by [20, 5] to accelerate the inference of LLMs. It utilizes a small model to generate the next $\gamma$ tokens and then uses the large model to verify the drafted tokens in parallel, which is summarized below:
@@
    Let $prefix$ be the input context, the small model samples $\gamma$ draft tokens $x_{1},\ldots,x_{\gamma}$ using greedy decoding based on the warped predicted conditional probability $\tilde{q}(x_{t}|x_{1:t-1},prefix))$ for $t=1,\ldots,\gamma$, where $\tilde{q}=\mathcal{T}\circ q$ and $q$ is the small model’s output distribution.
@@
    The large model verifies the draft tokens in parallel by computing the conditional probability $\tilde{p}(x_{t}|x_{1:t-1},prefix)$ for $t=1,\ldots,\gamma$.
@@
    Each draft token $x_{t}$ is accepted with probability $\min(1,\tilde{p}(x_{t})/\tilde{q}(x_{t}))$. The draft tokens before the first rejected token are kept as the decoding output. An additional token is sampled from a residual distribution as a correction to the first rejected token. Then the accepted tokens and the resampled token are appended to the context $prefix$ as the input to the next iteration.
@@
Because the large model verifies $\gamma$ tokens in parallel with one run, the time cost is smaller than calling it $\gamma$ times. Meanwhile, although the small model still runs in an autoregressive way, its inference speed is much faster than the large model. So speculative decoding can accelerate the inference process of LLMs. Additionally, it is proven that each token $x_{i}$ generated by speculative sampling follows the identical sampling distribution $\tilde{p}(x_{t}|x_{1:t-1},prefix)$ as greedy decoding [20]. However, since $\tilde{p}(x_{1:N}|prefix))\neq\prod_{t=1}^{N}\tilde{p}(x_{t}|x_{1:t-1},prefix)$, the effectiveness of speculative decoding also has room for improvement. Therefore, instead of having an identical sampling distributon as greedy decoding, our algorithm aims to achieve a better approximation to $\tilde{p}(x_{1:N}|prefix)$.
@@ 3 Methodology
The perplexity is a direct measurement of the effectiveness of a decoding algorithm. In practice, when the model is well-trained, a better perplexity usually leads to better downstream effectiveness. Although it is possible that improving perplexity does not improve the performance on downstream tasks, the problem is due to the model itself instead of the decoding algorithm. Therefore, the goal of this work is to design an algorithm that yields better output perplexity and better efficiency than greedy decoding and vanilla speculative decoding.
In this section, we first introduce multi-token joint greedy decoding (MJGD) to give a motivating example for why generating multiple tokens based on their joint likelihood can lead to better perplexity. Then we introduce multi-token joint speculative decoding (MJSD), which approximates and accelerates MJGD by exploiting a small model.
3.1 Multi-Token Joint Greedy Decoding
Let $M_{p}$ be the large target model, the goal is to generate a sequence of $N$ tokens with maximum likelihood $p(x_{1:N}|prefix)$. We now introduce a new decoding algorithm to improve greedy decoding in terms of perplexity.
@@ 3.1 Multi-Token Joint Greedy Decoding
Multi-Token Joint Greedy Decoding. Let $M_{p}$ be the large target model with distribution $p$. Different from single-token greedy decoding, multi-token joint greedy decoding (MJGD) generates the next $\gamma_{i}$ tokens at step $i$ based on their joint conditional probability $p(x_{t+1:t+\gamma_{i}}|x_{1:t})$ [^2] In the paper, we omit $prefix$ when there is no ambiguity., where $\gamma_{i}$ is an integer no less than 1 and $t=\sum_{i^{\prime}=1}^{i-1}{\gamma_{i^{\prime}}}$, i.e., the total tokens generated in the previous $i-1$ steps.
@@
Figure 1: Perplexity and rougeL score of the output when $\gamma_{i}=K$ for MJGD with OPT-125M and Llama-2-68M fined-tuned on ChatGPT-Prompts [26] dataset.
@@
Greedy decoding is a special case of MJGD where $\gamma_{i}=1$, $\forall i$. When $\gamma_{1}=N$, MJGD generates the optimal sequence by returning the sequence with the highest likelihood. So intuitively, output perplexity should improve as $\gamma_{i}$ increases. Additionally, generating $\gamma_{i}$ tokens simultaneously allows MJGD to return the optimal $\gamma_{i}$ tokens for step $i$. In contrast, vanilla greedy decoding selects each token without considering any future tokens. So MJGD is less prone to choosing local optima than greedy decoding.
@@
Empirical evidence supports our claim. Figure 1 shows the output perplexity and downstream ROUGE-L scores of MJGD with $\gamma_{i}$ set to a constant $K$, where $K=1,\ldots,5$. Notice that setting $K=1$ is equivalent to vanilla greedy decoding. We use beam decoding to approximate the $\arg\max$ sampling from the joint distribution $p(x_{t+1:t+K}|x+{1:t},prefix)$. We can see that when $K$ increases from 1 to 2, there is a significant improvement in perplexity. The perplexity continues to decrease when $K$ increases further. It confirms our claim that increasing $\gamma_{i}$ will increase the output perplexity. So $\prod_{i=1}^{n}\tilde{p}(x_{\Gamma_{i-1}+1:\Gamma_{i}}|prefix,x_{1:\Gamma_{i-1}})$ is a better approximation to $\tilde{p}(x_{1:N}|prefix)$ than $\prod_{t=1}^{N}\tilde{p}(x_{t}|prefix,x_{1:t-1})$. Here $\Gamma_{0}=0$, and $\Gamma_{i}=\sum_{j=1}^{i}\gamma_{i}$ for $i\geq 1$, and $n$ is the total number of steps of MJGD.
Moreover, the ROUGE-L score also improves with $K$, supporting our claim that better perplexity reflects enhanced performance in downstream tasks. This demonstrates that MJGD not only achieves lower perplexity but also yields higher quality text generation, making it a superior decoding strategy compared to vanilla greedy decoding [^3] Additional discussions comparing MJGD and greedy decoding are provided in the Appendix F..
3.2 Multi-Token Joint Speculative Decoding

### v2

# Optimized Multi-Token Joint Decoding with Auxiliary Model for LLM Inference

@@ Abstract
Large language models (LLMs) have achieved remarkable success across diverse tasks, yet their inference processes are hindered by substantial time and energy demands due to single-token generation at each decoding step. While previous methods such as speculative decoding mitigate these inefficiencies by producing multiple tokens per step, each token is still generated by its single-token distribution, thereby enhancing speed without improving output quality. In contrast, our work simultaneously enhances inference speed and improves the output effectiveness. We consider multi-token joint decoding (MTJD), which generates multiple tokens from their joint distribution at each iteration, theoretically reducing perplexity and enhancing task performance. However, MTJD suffers from the high cost of sampling from the joint distribution of multiple tokens. Inspired by speculative decoding, we introduce multi-token assisted decoding (MTAD), a novel framework designed to accelerate MTJD. MTAD leverages a smaller auxiliary model to approximate the joint distribution of a larger model, incorporating a verification mechanism that not only ensures the accuracy of this approximation, but also improves the decoding efficiency over conventional speculative decoding. Theoretically, we demonstrate that MTAD closely approximates exact MTJD with bounded error. Empirical evaluations using Llama-2 and OPT models ranging from 13B to 70B parameters across various tasks reveal that MTAD reduces perplexity by 21.2% and improves downstream performance compared to standard single-token sampling. Furthermore, MTAD achieves a 1.42$\times$ speed-up and consumes 1.54$\times$ less energy than conventional speculative decoding methods. These results highlight MTAD’s ability to make multi-token joint decoding both effective and efficient, promoting more sustainable and high-performance deployment of LLMs.
@@ 1 Introduction
Large Language Models (LLMs) such as GPT-4 and Llama-2 (Touvron et al., 2023) have demonstrated extraordinary capabilities across a wide range of tasks (Brown et al., 2020; Chowdhery et al., 2023; Thoppilan et al., 2022; Touvron et al., 2023). Despite their impressive performance, the deployment of LLMs is often constrained by substantial inference costs in terms of time and energy. This inefficiency primarily stems from the autoregressive nature of these models, where generating a sequence of $K$ tokens requires $K$ separate model calls. Each call involves loading large weight matrices and intermediate results from GPU global memory to computing units, leading to repeated memory accesses and limited hardware utilization (Samsi et al., 2023; Leviathan et al., 2023).
@@
To tackle this challenge, researchers have delved into non-autoregressive decoding approaches. Early methods (Ghazvininejad et al., 2019; Gu et al., 2017; Guo et al., 2020) aimed at reducing inference latency by concurrently generating multiple tokens. But these methods usually require task-dependent techniques and information to match the performance of autoregressive decoding (Kim et al., 2023; Xiao et al., 2023). More recently, speculative decoding has emerged (Leviathan et al., 2023; Chen et al., 2023; Kim et al., 2023; Sun et al., 2023), exploiting the observation that most of the small model’s prediction aligns well with that of a large model. It leverages a smaller auxiliary model to draft a few future tokens autoregressively, which are subsequently validated in parallel by the larger model. As the smaller model operates significantly faster and parallel token verification incurs a similar time cost as generating a single token, speculative decoding achieves an overall speed-up of $1$-2×2$\times$2 ×. Despite gains in speed, these methods still generate each token based on its single-token probability. Consequently, it does not enhance the effectiveness of the generated sequences.
In this work, we aim to go beyond the conventional trade-off between efficiency and effectiveness by introducing multi-token joint decoding (MTJD). Unlike traditional approaches, MTJD produces multiple tokens from their joint distribution at each decoding step. Theoretically, we show this joint generation can lead to lower perplexity and hence improved task performance. However, directly sampling from the joint distribution of multiple tokens poses significant computational challenges, rendering MTJD impractical.
Inspired by speculative decoding, we propose multi-token assisted decoding (MTAD), a novel framework designed to approximate and accelerate MTJD. MTAD employs a smaller auxiliary model to estimate the joint distribution of a larger model, significantly reducing computational demands. To ensure the accuracy of this approximation, MTAD incorporates a verification mechanism that not only guarantees the accuracy of the draft tokens but also enhances efficiency beyond conventional speculative decoding by maximizing the number of accepted tokens per iteration. We provide both theoretical and empirical analyses to demonstrate that MTAD improves perplexity and downstream performance. Meanwhile, it achieves significant reductions in energy and time usage compared to existing decoding strategies.
Our contributions are as follows:
  1. 1.
    We introduce multi-token joint decoding (MTJD), a multi-token joint decoding approach that theoretically reduces perplexity by generating tokens from their joint distribution.
  2. 2.
    We develop multi-token assisted decoding (MTAD), an efficient approximation of MTJD with bounded error that leverages a smaller model for distribution approximation.
  3. 3.
    We analyze the energy consumption of LLM inference. To our knowledge, we are the first to give theoretical and empirical evidence that, despite that MTAD and other speculative decoding algorithms increase the number of FLOPs needed during LLM inference, they reduce the overall energy consumption by reducing the overhead induced by accessing GPU global memory.
  4. 4.
    We conducted comprehensive evaluations with Llama-2 and OPT models (ranging from 13B to 70B parameters) across various tasks, demonstrating that MTAD enhances perplexity by 21.2% and improves downstream effectiveness compared to standard single-token sampling, while also achieving a 1.42$\times$ speed-up and reducing energy consumption by 1.54$\times$ compared to conventional speculative decoding methods.
These advancements position MTAD as a robust solution for making multi-token joint decoding both effective and efficient, thereby facilitating more sustainable and high-performance deployment of large-scale language models. Our code is publicly available [^1] https://anonymous.4open.science/r/LLMSpeculativeSampling-EE52.
@@ 2.1 Decodings of LLMs
Decoding and Perplexity.
Let $p$ denote the distribution defined by LLM model $M_{p}$. Given an input context $input$, a decoding algorithm generates a sequence of $N$ tokens whose likelihood is denoted as $p(x_{1:N}|input)$. The likelihood of the sequence is directly linked to perplexity of the sequence, which is the exponentiated average negative log-likelihood of all tokens. Based on autoregressive decomposition $p(x_{1:N}|input)=\prod_{t=1}^{N}p(x_{t}|x_{1:t-1},input)$ [^2] In the paper, we omit $input$ when there is no ambiguity., the perplexity is defined as:
@@
Perplexity serves as a direct metric for assessing the effectiveness of a decoding algorithm. In practice, when a model is well-trained, lower perplexity often correlates with improved downstream performance. For example, beam sampling aims to return output with lower perplexity and is proven to have better downstream performance in general (Shi et al., 2024).
To further demonstrate the relationship between perplexity and downstream performance, we evaluate GPT-3.5-turbo on the spider (Yu et al., 2018) dataset. Using a temperature of 2, the model generated 10 outputs for each input. We measured the average perplexities and execution accuracies for the outputs with the highest, lowest, and median (the 5-th lowest) perplexity. As shown in Table 1, lower perplexity correlates with improved downstream performance, even in one of today’s largest models.
Table 1: Relationship between perplexity and execution accuracy (EA, higher the better) for GPT-3.5-turbo.
Output Avg. PPL $\downarrow$ EA (%) $\uparrow$
Highest PPL 4.13 33
5-th Lowest PPL 1.40 58
Lowest PPL 1.07 62
@@
Multinomial Sampling.
Multinomial sampling, also known as standarized sampling or single-token sampling, samples the next token $x_{t}$ based on $\mathcal{T}\circ p(\cdot|x_{1:t-1},input)$, where $\mathcal{T}$ is a warping operation applied to enhance the high probability region. Some common warping operations include top-k warping, which limits the selection to the top k tokens, and top-p warping, where tokens are sampled from the smallest possible subset of the vocabulary whose cumulative probability mass exceeds a specified threshold. The deterministic version of multinomial sampling (i.e., greedy decoding) is a special case when $k=1$.
Beam Sampling.
Beam sampling aims to improve output perplexity over multinomial sampling. For each position $t$ ($1\leq t\leq N$), it maintains $W>1$ candidate sequences, which are also called beams. Assume we have already kept the $W$ sequences $\mathcal{I}_{t-1}=\{x_{1:t-1}^{(1)},\ldots,x_{1:t-1}^{(W)}\}$ at position $t-1$, $W$ sequences with length $t$ are then sampled from $\mathcal{T}\circ p_{beam}$, where $p_{beam}:\mathcal{I}_{t-1}\times V\rightarrow[0,1]$ is the beam sampling probability:
$p_{beam}(x_{1:t-1}^{(i)},x_{t})=\frac{p(x_{1:t-1}^{(i)},x_{t}|input)}{\sum_{x_{1:t-1}^{(j)},x^{\prime}_{t}\in\mathcal{I}_{t-1}\times V}p(x_{1:t-1}^{(j)},x_{t}^{\prime}|input)}$ (2)
Notice that $p(x_{1:t-1}^{(i)},x_{t}|input)=p(x_{t}|x_{1:t-1}^{(i)},input)\cdot p(x_{1:t-1}^{(i)}|input)$. In practice, beam sampling stores the likelihood $p(x_{1:t-1}^{(i)}|input)$ for each beam, and the computation complexity of $p_{beam}$ is $O(W\cdot|V|)$. In deterministic beam sampling, the top $W$ sequences with the highest likelihood $p_{beam}(x_{1:t})$ will be kept.
@@ 2.2 Vanilla Speculative Decoding
Besides effectiveness, speculative decoding is proposed by (Leviathan et al., 2023; Chen et al., 2023) to accelerate the inference of LLMs. It utilizes a small model to generate the next $\gamma$ tokens and then uses the large model to verify the drafted tokens in parallel, which is summarized below:
@@
    Let $input$ be the input context, the small model samples $\gamma$ draft tokens $x_{1},\ldots,x_{\gamma}$ using multinomial sampling based on $\tilde{q}(x_{t}|x_{1:t-1},input))$ for $t=1,\ldots,\gamma$, where $\tilde{q}=\mathcal{T}\circ q$ and $q$ is the small model’s output distribution.
@@
    The large model verifies the draft tokens in parallel by computing the conditional probability $\tilde{p}(x_{t}|x_{1:t-1},input)$ for $t=1,\ldots,\gamma$.
@@
    Each draft token $x_{t}$ is accepted with probability $\min(1,\tilde{p}(x_{t})/\tilde{q}(x_{t}))$. The draft tokens before the first rejected token are kept as the decoding output. An additional token is sampled from a residual distribution as a correction to the first rejected token. Then the accepted tokens and the resampled token are appended to the context $input$ as the input to the next iteration.
@@
Because the large model verifies $\gamma$ tokens in parallel with one run, the time cost is smaller than calling it $\gamma$ times. Meanwhile, although the small model still runs in an autoregressive way, its inference speed is much faster than the large model. As a result, speculative decoding achieves a speedup of 1–2× compared to multinomial sampling while maintaining an identical sampling distribution.
@@ 3 Methodology
As discussed in Section 2, the goal of this work is to design an algorithm that yields lower perplexity and better efficiency than multinomial sampling and vanilla speculative decoding. In this section, we first introduce multi-token joint decoding (MTJD), which generates multiple tokens based on their joint likelihood. We prove it can yield lower perplexity. Then we introduce multi-token assisted decoding (MTAD), which approximates and accelerates MTJD by exploiting an auxiliary model.
3.1 Multi-Token Joint Decoding
We first introduce a new decoding algorithm to improve multinomial sampling in terms of perplexity.
@@ 3.1 Multi-Token Joint Decoding
Multi-Token Joint Decoding. Let $M_{p}$ be the large target model with distribution $p$. Different from single-token multinomial sampling, multi-token joint decoding (MTJD) generates the next $\gamma_{i}$ tokens at step $i$ based on their joint conditional probability $p(x_{t+1:t+\gamma_{i}}|x_{1:t})$, where $\gamma_{i}$ is an integer no less than 1 and $t=\sum_{i^{\prime}=1}^{i-1}{\gamma_{i^{\prime}}}$, i.e., the total tokens generated in the previous $i-1$ steps.
@@
Figure 1: Perplexity and Rouge-L score of the output when $\gamma_{i}=K$ for MTJD with OPT-125M and Llama-2-68M fine-tuned on ChatGPT-Prompts (Rashad, 2023) dataset.
@@
Multinomial sampling is a special case of MTJD where $\gamma_{i}=1$, $\forall i$. When $\gamma_{1}=N$, MTJD generates the sequence directly based on their joint likelihood. So intuitively, output perplexity should improve as $\gamma_{i}$ increases. Besides, generating $\gamma_{i}$ tokens simultaneously allows MTJD to consider their interactions. In contrast, multinomial sampling selects each token without considering any future tokens. So MTJD is less prone to choosing local optima.
Theorem 3.2 shows the limit of perplexity of MTJD when $N$ approaches infinity. The proofs are included in the Appendix A.
Theorem 3.2.
Assume at the $i$-th ($i=1,\ldots,N$) iteration, MTJD generates $\gamma_{i}$ tokens. Let $\Gamma_{i}$ denote the total number of tokens generated at the first $i$ iterations. Let $x_{1:\Gamma_{N}}$ denote the generated tokens. When $N\rightarrow\infty$
$PPL_{p}(x_{1:\Gamma_{N}})\rightarrow\exp\left(-\frac{1}{\bar{\gamma}}\mathbb{E}_{\gamma}L_{p}(\gamma,\tilde{p})\right)$ (3)
where $\bar{\gamma}$ is the expected number of $\gamma_{i}$, $\tilde{p}=\mathcal{T}\circ p$ represents how we sample the next $\gamma_{i}$ tokens from $p$ (e.g., in deterministic sampling, $\tilde{p}=\arg\max\circ p$ always returns the tokens with the highest joint likelihood), and $L_{p}(\gamma,\tilde{p})$ is the expected log-likelihood of the $\gamma$ tokens sampled from $\tilde{p}$:
$L_{p}(\gamma,\tilde{p})=\mathbb{E}_{x_{1:t}\in\mathcal{X}}\sum_{x_{t+1:t+\gamma}}\tilde{p}(x_{t+1:t+\gamma}|x_{1:t})\log p(x_{t+1:t+\gamma}|x_{1:t})$ (4)
Here $\mathcal{X}$ is the space of all possible inputs.
Corollary 3.3.
Based on Theorem 3.2, we can show that when $N\rightarrow\infty$, greedy MTJD (i.e., top-1 MTJD sampling) has lower perplexity than greedy decoding (top-1 single-token sampling).
@@
Empirical evidence supports our claim. We fine-tune both a Llama and an OPT model on the ChatGPT-Prompts dataset and evaluate the output perplexity and Rouge-L scores with example outputs. Figure 1 shows the output perplexity and Rouge-L scores of MTJD with $\gamma_{i}$ set to a constant $K$, where $K=1,\ldots,5$. Notice that setting $K=1$ is equivalent to multinomial sampling. We use beam sampling to approximate the $\arg\max$ sampling from the joint distribution $p(x_{t+1:t+K}|x+{1:t},input)$. We can see that the perplexity keeps dropping when $K$ increases. It confirms our claim that increasing $\gamma_{i}$ will increase the output perplexity. Moreover, the Rouge-L score also improves with $K$, supporting our claim that better perplexity reflects enhanced performance in downstream tasks.
3.2 Multi-Token Assisted Decoding